00:38
Run 25 started - Beam on target
Beam energy = 1.5 MeV.
Horizontal beam position = 7.0 mm, vertical beam position = 0.0 mm.
Target density ~ 5e11 at/cm2 (Still Fluctuating, after temperature increase)
Beam intensity at inj ~ 2.5e7
Detectors are positioned IN. Left arm = 62.71, right arm = 64.60 mm
Histograms and stats zero'd, monitoring code restarted.
Beam in the ring ~ 5 minutes.
Pressure in YR09 E target ~ 5.2e-11 mbar.
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
01:30 (Attachments 1 - 4)
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
The target density seems very unstable: fluctuations from 1e11 to 1e12 atoms/cm2 within 5min
02:37 (Attachments 5 - 8)
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
The target density is around 5e11 atoms/cm2 with small adjustments
03:32 (Attachments 9 - 12)
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
The target density is in average 5e11 atoms/cm2 with some small adjustments. Max values up to 3e12 atoms/cm2
One of the injections at 03:43 was not ok and we lost beam for some min. Target density 1e12 atoms/cm2. On 04:00 the beam was back again without any operator intervention.
04:35 (Attachments 13 - 16)
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
The target density is in average 1e12 atoms/cm2.
05:43 (Attachments 17 - 20)
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
The target density decreased to 2e11 atoms/cm2 in the last 40 min
06:43 (Attachments 21 - 24)
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
The target density is fluctuating from 1e11 to 1e12 atoms/cm without any change....
07:53 (Attachments 25 - 28)
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
The target density is fluctuating from 1e11 to 1e12 atoms/cm without any change....
08:48 (Attachments 29 - 32)
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
The target density is fluctuating from 1e11 to 1e12 atoms/cm without any change....
09:40 --> Run 25 stopped.
|